# SADMoE: Exploiting Activation Sparsity with Dynamic-$k$ Gating

### v2

@@ 4.1 Language modeling
Figure 6: Test accuracy of BERT models on CARER dataset at different compute levels. SADMoE displays significant improvements over MoEfication.
@@
We evaluate our method on language modeling using GPT-2-base Radford et al. (2019) architecture by comparing it to MoEfication. We initialize the models from a publicly available OpenAI checkpoint pre-trained on a closed-source WebText dataset, and use OpenWebText (Gokaslan and Cohen, 2019) in all of our experiments. We train our model with sparsity enforcement loss on OpenWebText for 1000 steps, as we find out that it is sufficient for the model to recover the original loss value. We report the results in this Section without the MHA projection replacement step, as we find that GPT is especially sensitive to changes in attention layers, which always result in visible loss degradation.
Test losses for both methods are presented for different levels of compute budgets in Figure 5. Our method outperforms the baseline at every computational budget, owing its performance to the increased activation sparsity of the model and its ability to dynamically allocate the needed compute. Notably, the loss of SADMoE plateaus for higher budget levels, while the baseline displays consistently worse results whenever we lower the computational budget.
4.2 Text classification
We evaluate our method with BERT-base (Devlin et al., 2018) on the CARER dataset Saravia et al. (2018) that contains text samples categorized into 6 different emotion categories. We compare the accuracy versus FLOPs trade-off for SADMoE and MoEfication. For broader comparison, we also include a state-of-the-art early-exit method Zero-time Waste (ZTW) Wójcik et al. (2023). However, we should note that early-exit methods are orthogonal to MoE methods. The results are shown in Figure 6.
The performance of MoEfication gradually deteriorates and completely collapses when the number of executed experts approaches zero. In comparison, SADMoE maintains the original performance for a wide range of computational budgets, and the performance drop starts at a significantly lower budget. While early-exiting performs well for the lowest budgets, it obtains worse results than SADMoE at medium budgets and also suffers from gradual performance decline.
4.3 Image classification
Figure 7: ImageNet-1k FLOPs vs testset accuracy results for SADMoE, MoEfication and ZTW applied to ViT-B. Our method clearly outperforms the baselines on image classification due to elevated activation sparsity levels and dynamic-$k$ expert selection.
Vision Transfomers (Dosovitskiy et al., 2020) are becoming one of the most popular architectures in computer vision. Since our method is applicable to any Transformer model, we evaluate our method on the popular ImageNet-1k (Russakovsky et al., 2015) dataset. We use a pre-trained ViT-B checkpoint from the torchvision library [^2] https://pytorch.org/vision/stable/models.html as the base model and compare SADMoE with MoEfication and ZTW in terms of computational cost versus accuracy trade-off. Our results, presented in Figure 7, demonstrate the significant gains from applying our method over MoEfication, which emphasizes the role of induced activation sparsity and dynamic-$k$ expert selection rule.
@@ 5 Analysis
In this section, we present the results of experiments that provide insights into the performance of our method. Unless stated otherwise, we follow the experimental setup for models and datasets from the previous section.
Figure 8: Performance of ablated variants of SADMoE and MoEfication with BERT-base on the CARER dataset. We label the components of our method as S: sparsity enforcement, RR: regression routing, DynK: dynamic-$k$, and A: extension to attention projections. Each of the components of SADMoE meaningfully improves performance.
@@ 5.1 Ablation study
Since our method consists of several steps, the positive impact of each one of them may not be evident. To show the significance of every component, we perform an ablation study by incrementally adding each component to the baseline method. We take a BERT-base model and evaluate the ablated variants on CARER dataset in the same setting as described in Section 4.2. The results of this experiment are presented in Figure 8. As expected, each ablated version of the method improves upon the previous one. The sparsity enforcement phase leads to enhanced performance compared to plain MoEfication, as sparser activations enable better reconstruction of original output with fewer experts. Alternative router training objective and dynamic-$k$ expert assignment further improve the results, but – as the method only operates on the FFN layer – the computational cost cannot go below the cost of the remaining part of the model. Extending SADMoE to MHA projection layers allows our method to further reduce the computational cost, and the resulting full method retains the accuracy of the original model at around twice fewer FLOPs than MoEfication.
5.2 Sparsification and reliance on activation function
Figure 9: Comparison of our method applied on GPT-2 models with GELU and ReLU activation functions, with and without the activation sparsity enforcement phase. The sparsification enables the standard GELU model to perform similarly to the ReLU-based one, while the non-sparsified GELU model is inferior to its ReLU counterpart.
Activation sparsity works focus their analysis on networks with ReLU, as activation functions commonly used in Transformer models (such as GELU or SiLU) do not output exact zeros for inputs below zero. Although activations with near-zero values might have a negligible impact on the final output, assessing their number is non-trivial. They also make it hard to apply activation sparsity enforcement schemes, such as $\ell^{1}$ penalty, that mostly assume the ReLU activation function. To get around this limitation, the works on activation sparsity usually change the activation function of the trained model to ReLU and continue pre-training the modified model until it reaches the performance of the original model  (Zhang et al., 2022; Mirzadeh et al., 2023).
Figure 10: Effect of expert division granularity on our method. In general, smaller experts display favorable performance as they allow for larger computational savings.
We hypothesize that the dependence on ReLU is not necessary and that the models with many near-zero activations can effectively function in a very similar way to standard activation-sparse ReLU-based models. To evaluate this hypothesis, we extend the sparsity enforcement scheme to the GELU activation function by penalizing the model for pre-activation values larger than a certain threshold. We first transform the pre-activation values:
$z^{\prime}=\max(0,z-d),$ (5)
@@ 5.2 Sparsification and reliance on activation function
where $z$ is the pre-activation value and $d$ is the displacement hyperparameter. Then, we apply the loss from Equation 1 on $z^{\prime}$. The purpose of this transformation is to penalize the model only for pre-activation values larger than $d$. We empirically find that $d=-10$ works well with GELU activation, as the output of GELU activation function below this value is near-zero. Following the methodology from Section 4.1, we train ReLU-based and GELU-based GPT-2 with and without sparsity enforcement loss and show the results of our experiment in Figure 9. SADMoE with a sparsified GELU-based model performs similarly to a sparsified ReLU-based model, while the performance of the non-sparsified GELU-based variant rapidly collapses for $\tau>0$. Within ReLU-based models, the sparsification still enhances the performance of SADMoE, but the improvements are less drastic, and the behavior of our method does not significantly change as in the case of GELU. This shows sufficient activation sparsity enforcement relieves the model from the dependence on ReLU.
Figure 11: Per-layer distribution of the number of executed experts in SADMoE trained on the CARER with different $\tau$ thresholds for a standard, non-sparsified model (top row) and a sparsified model (bottom row). The high variability of that number explains the computational gains from using dynamic-$k$.
@@ 5.3 Impact of expert granularity
A crucial hyperparameter in SADMoE is the selection of expert size. Smaller experts may allow for a more granular selection of executed neurons, thus likely resulting in a lower computational cost. However, decreasing the expert size increases the number of experts, which translates to a larger router, potentially negating any computational gains. To study the impact of this hyperparameter on our method, we evaluate SADMoE on GPT-2 with experts of different sizes, and show the results in Figure 10.
We observe that our method generally performs better with smaller experts. Those results differ from the results presented in  (Zhang et al., 2022), where the expert size is significantly higher. The positive correlation between granularity and performance can be explained by the increased levels of activation sparsity in our model, which requires significantly fewer activated neurons (experts) to correctly function. As expected, the performance decreases for the extreme choice of expert size equal to 1 due to significantly larger routing costs. We include additional results for expert granularity with GELU activation in Appendix D.
5.4 Expert selection patterns
To explore the variability in SADMoE expert usage patterns, we investigate the distribution of the number of executed experts for different values of the $\tau$ threshold, with and without the sparsification phase. In Figure 11, we show the histograms of the numbers of activated experts for each layer FFN of the BERT-base trained on the CARER dataset. As expected, the model with enforced activation sparsity requires a smaller number of experts for each given threshold threshold. Deeper model layers generally utilize more experts, which suggests that the features at those levels are more complex. Interestingly, even for high values of $\tau$, the sparsified model does not always use all available experts. Both base and sparsified models also exhibit significant variance in the number of activated neurons across different layers, which justifies the dynamic-$k$ selection and indicates that computational adaptability mechanisms are crucial for efficient inference in Transformer-based models.
@@ 6 Conclusion
We introduce Sparsified Activation Dynamic-$k$ Mixture-of-Experts (SADMoE), a novel approach that induces activation sparsity to improve the efficiency of Transformer-based models by converting their layers to Mixture-of-Experts (MoE). We demonstrate the interplay between the activation sparsity of dense models and the efficiency of converted MoE. Moreover, we introduce regression-based router training and dynamic-$k$ routing, which enable our method to efficiently utilize the induced sparsity. Finally, we show how dense-to-sparse-MoE conversion approaches can be extended to standalone dense layers. Our approach is compatible with existing Transformer architectures like GPT, BERT, and ViT, and significantly improves upon the existing MoE conversion schemes. Our findings contribute to the ongoing efforts to make Transformer models more efficient and accessible for a wider range of applications, especially in resource-constrained environments.
Limitations
While SADMoE displays promising results in reducing the computational cost of inference in Transformer models, several limitations should be acknowledged. Our proposed sparsity enforcement and router training phases require additional training time. This overhead, while small, should be considered when evaluating the overall efficiency gains of our approach. Moreover, we demonstrate improved performance over existing approaches on common NLP and vision tasks, but the scope of our experiments is restricted due to limited access to computational resources. Further research is needed to address the aforementioned limitations and to explore its applicability to extremely large models. We focus primarily on fundamental machine learning research and we do not see any specific risks or ethical issues associated with our method. Nevertheless, we recognize the potential for misuse of machine learning technology and advocate for responsible AI practices to mitigate such risks.
@@ Acknowledgments
We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, PCSS, CI TASK, WCSS) for providing computer facilities and support within computational grants no. PLG/2023/016393 and PLG/2023/016321. Filip Szatkowski is supported by National Centre of Science (NCP, Poland) Grant No. 2022/45/B/ST6/02817. Bartosz Wójcik is supported by National Science Centre, Poland, Grant No. 2023/49/N/ST6/02513.
Appendix A Difference between router training in SADMoE and MoEfication
Our router training procedure might seem similar to the one proposed in MoEfication (Zhang et al., 2022), but the source code of the method provided by the authors [^3] MoEfication source code for router training is publicly available at: https://github.com/thunlp/MoEfication/blob/c50bb850307a36f8a0add6123f56ba309a156d13/moefication/utils.py#L188-L260 contains a different routing objective than the one reported in the paper. While the paper describes their router training objective as a prediction of the sum of ReLU activation values in each expert, the source code uses prediction labels created from the sum of the activations in the intermediate layer divided by the maximum value in the entire batch and minimize the binary cross-entropy loss. Assuming that $a_{k,j}$ is the activation vector in the hidden layer of expert $j$ for sample $k$, the label generation for their router can be expressed as:

### v4

@@ 4.3 Language modeling
We evaluate our method on language modeling and compare it with MoEfication using GPT-2-base [31] and Gemma-2B [42]. We initialize GPT-2 models from a publicly available OpenAI checkpoint pre-trained on a closed-source WebText dataset and use OpenWebText [12] in all of our experiments. For Gemma-2B, we also start from the publicly available pretrained model and evaluate its language capabilities on the C4 dataset [32] after finetuning. For both models, we use around 1B tokens for the finetuning phase (less than 1% of the cost of original pretraining) and 8-16M tokens for router training. We report the results in this section without the MHA projection replacement, as this task is highly sensitive to changes in attention layers, leading to noticeable loss degradation. For more training details, see Section J.3
We present test losses for D2DMoE and MoEfication at different compute budgets for GPT-2-base and Gemma-2B in Figures 4(c) and 4(d) respectively. Our method outperforms the baseline at every computational budget. The loss of D2DMoE plateaus for higher budget levels, while the baseline displays consistently worse results whenever we lower the computational budget. Notably, for the larger Gemma-2B model our method performs well for most compute budgets, while the performance of MoEfication collapses. The failure of MoEfication can be explained by the emergence of massive activations in large models [40], which makes it unable to learn reliable routing, as we analyze in more detail in Appendix E.
We also provide a downstream evaluation of our Gemma models on the BoolQ dataset. We take the base model, which achieves 68.40% zero-shot evaluation accuracy, and convert it to MoE with D2DMoE and MoEfication. In Table 1, we report the relative accuracy of the models at different compute budgets. Our method largely retains the performance across multiple compute budgets, while the performance of MoEfication decreases significantly. This shows that the loss-vs-FLOPs results for D2DMoE and MoEfication directly translate to downstream performance on language tasks.
Table 1: Relative downstream performance of D2DMoE and MoEfication on BoolQ dataset. Our method only starts to degrade at around 70% compute budget, while MoEfication gradually decreases.
Compute budget 100% 90% 80% 70% 60% 50% 25% 10%
MoEfication 100% 92.24% 92.19% 92.15% 88.79% 75.40% 86.70% 77.53%
D2DMoE 100% 99.68% 99.37% 98.69% 97.60% 94.34% 92.75% 90.89%
4.4 Execution latency
Figure 5: Single D2DMoE layer execution wall-clock time.
For any model acceleration method to be practically useful, it must reduce end-to-end inference execution time on modern GPU hardware. To achieve this, we implement the forward pass of our MoE-layer in the Triton intermediate language [43], and employ several optimizations for our implementation, including an efficient memory access pattern, kernel fusion, and configuration auto-tuning. As suggested by Tan et al. [41], our implementation also avoids unnecessary copies when grouping tokens.
We verify the performance of our implementation for a single D2DMoE layer ($24$ experts with expert dimensionality $128$) layer in isolation by comparing it with the corresponding MLP module (inner dimensionality $3072$) on an NVIDIA A100 GPU. We fill a tensor of size $[256\times 197\times 768]$ (batch size, sequence length, and hidden dimension, respectively) filled with Gaussian noise and use it as input to both modules. The gating network of D2DMoE is included in measurements, but the decisions are overridden with samples from a Bernoulli distribution, and we control how many experts are executed on average by changing the Bernoulli probability. The results, presented in Figure 5, show that our implementation scales linearly with the number of executed experts, and has negligible overhead. Our method can be almost three times as fast as standard MLP while preserving 99% of the original accuracy. In Appendix C we provide additional wall-clock measurement results along with a more detailed description of our implementation.
4.5 Compatibility with model compression techniques
Figure 6: D2DMoE applied to models pruned with CoFi.
Figure 7: D2DMoE applied to quantized models.
To accelerate inference D2DMoE leverages input-dependent activation sparsity, a property inherent to almost every Transformer model. However, interaction between D2DMoE and other popular network acceleration techniques, such as pruning [16] or quantization [13, 28], is unclear. We evaluate D2DMoE in combination with such techniques to demonstrate their complementarity.
First, we evaluate D2DMoE applied on top of networks pruned with CoFi, a structured pruning technique introduced by Xia et al. [50]. CoFi removes redundant neurons, attention heads, and sublayers to achieve the desired sparsity ratio, and then subsequently fine-tunes the reduced network. We first prune the base model with CoFi to the desired sparsity level, apply D2DMoE to it, and then evaluate both models on QNLI [48]. In Figure 7, we show that D2DMoE successfully accelerates inference even on networks pruned to high sparsity levels.
In  Figure 7, we also investigate the applicability of D2DMoE to quantized models using dynamic post-training quantization from PyTorch [^2] https://pytorch.org/tutorials/recipes/recipes/dynamic_quantization.html on BERT trained on the CARER dataset. Our method is robust to 8- and 16-bit quantization and exhibits only slight variations in performance after quantization. As FLOPs do not take bit width into account, we show quantized models in the same FLOPs range as the original model. In Appendix C, we also present wall-clock time measurements for quantized D2DMoE.
@@ 5 Analysis
In this section, we present in detail additional experiments that provide insights into the performance of our method. Additionally, in Appendix E we analyze the performance of MoEfication with Gemma, in Appendix F we provide the results of router architecture analysis, in Appendix H we conduct experiments corresponding to the ones in Section 5.5 with GELU function, and in Appendix I we show additional visualizations for expert activation patterns.
5.1 Expert selection patterns
(a) Compute along the model depth
(b) Computational load maps for ImageNet-1k sample images
Figure 8: D2DMoE allows for a dynamical allocation of computation for each layer and each input independently. a) Per-layer distribution of the number of executed experts on CARER dataset in D2DMoE with $\tau=0.01$ for a standard model (top) and a sparsified model (bottom). Sparsification leads to a significantly lower number of selected experts. b) Computational load maps of selected ImageNet-1k samples for our converted ViT-B model with $\tau=0.0025$. D2DMoE allocates its computational budget to semantically important regions of the input.
The dynamic-$k$ rule introduces variability in the allocation of the computational budget along the model depth. To explore its scale, we investigate the distribution of the number of executed experts, with and without the activation sparsification phase. In Figure 8(a), we show the histograms of the number of activated experts for each FFN layer of the BERT-base model trained on the CARER dataset (additional results are available in the appendix in Appendix I). As expected, the model with enforced activation sparsity requires fewer experts for a given threshold. Both base and sparsified models exhibit significant variance in the number of activated neurons across different layers, which justifies the dynamic-$k$ selection and indicates that computational adaptability mechanisms are crucial for efficient inference in Transformer-based models.
D2DMoE also allows the model to allocate different computational resources to various layers. We expect the model to allocate more compute to tokens containing information relevant to the current task. Since each token position in a ViT model corresponds to a separate and non-overlapping part of the input image, we can easily plot a heatmap to indicate the regions of the image where the model spends its computational budget. In Figure 8(b) we present such an analysis for our converted ViT-B model. As expected, the dynamic-$k$ routing enables the model to minimize the computational effort spent on regions that contain insignificant information.
@@ 5.2 Ablation study
Since our method consists of several steps, the positive impact of each one of them may not be evident. To show the significance of every component, we perform an ablation study by incrementally adding each component to the baseline method. We take a BERT-base model and evaluate the ablated variants in the same setting as described in Section 4.2. The results of this experiment are presented in Figure 9(a). As expected, each ablated version of the method improves upon the previous one. The sparsity enforcement phase leads to enhanced performance compared to plain MoEfication. Alternative router training objective and dynamic-$k$ expert assignment further improve the results, but – as the method only operates on the FFN layer – the computational cost cannot go below the cost of the remaining part of the model. Extending D2DMoE to MHA projection layers allows our method to reduce the computational cost further, and the resulting full method retains the accuracy of the original model at around twice fewer FLOPs than MoEfication.
5.3 Base model activation sparsity
To justify our proposed activation sparsity phase, we investigate the impact of the activation sparsity of the base dense model on the performance MoE obtained with our method. We conduct a study similar to the one presented in Figure 2(a): we train multiple base models with different activation sparsity enforcement loss weights $\alpha$ and convert them to Mixture-of-Experts models with our method.
The results, shown in Figure 9(b), highlight the positive correlation between the activation sparsity and the performance of the converted MoE, as higher sparsity in the base model always translates to better performance for D2DMoE. This is consistent with results previously observed for MoEfication. However, our method achieves better results for every base model in all cases, proving that regression routing and dynamic-$k$ selection better utilize the induced sparsity.
(a) Phases ablation (BERT-base)
(b) Sparsification (GPT2-base)
(c) ReLU vs GELU (GPT2-base)
(d) Expert granularity (GPT2-base)
Figure 9: Analysis experiments with D2DMoE. (a) Impact of different phases of our method. Each phase improves upon the baseline. (b) Sparsification improves the cost-accuracy trade-off of the final D2DMoE model. (c) Sparsification allows us to apply our method to GELU-based model without significant drops in performance. (d) Smaller experts display favorable performance and allow for larger computational savings.
5.4 Sparsification and reliance on the activation function
Activation sparsity works focus their analysis on networks with ReLU activation, as other functions (such as GELU or SiLU) do not guarantee exact sparsity. When analyzing non-ReLU models, such works require fine-tuning with the activation function changed to ReLU (relufication[52, 27], which limits their practical applicability. We hypothesize that relufication is not necessary and the models with many near-zero activations effectively function similarly to standard ReLU-based models. To evaluate this hypothesis, we extend the sparsity enforcement scheme to the commonly used GELU activation by penalizing the model for pre-activation values larger than a certain threshold. We first transform the pre-activation values as $z^{\prime}=\max(0,z-d)$, where $z$ is the pre-activation value and $d$ is a displacement hyperparameter. Then, we apply the loss from Equation 1 on $z^{\prime}$. This transformation penalizes only pre-activation values larger than $d$, and as a result, the model learns to produce values that effectively become negligible post-activation. We empirically find that $d=-10$ works well with GELU as the output below this value is near zero.
@@ 5.4 Sparsification and reliance on the activation function
To validate our hypothesis, we follow the methodology from Section 4.3 and we train ReLU- and GELU-based GPT-2 with and without sparsity enforcement loss. We show the results in Figure 9(c). D2DMoE with a sparsified GELU-based model performs similarly to a sparsified ReLU-based model, while the performance of the non-sparsified GELU-based variant collapses. Within ReLU-based models, the sparsification still enhances the performance of D2DMoE, but the improvements are less drastic, and the behavior of our method does not significantly change as in the case of GELU. This shows sufficient activation sparsity enforcement relieves the model from the dependence on ReLU.
@@ 5.5 Impact of expert granularity
A crucial hyperparameter in D2DMoE is the selection of expert size. Smaller experts may allow a more granular selection of executed neurons, likely resulting in a lower computational cost. However, decreasing the expert size increases the number of experts, which translates to a larger router, potentially negating any computational gains. To study the impact of this hyperparameter on our method, we evaluate D2DMoE on GPT-2 with different expert sizes, and show the results in Figure 9(d).
We observe that our method generally performs better with smaller experts. Those results differ from the ones presented in  [52], where the expert size is significantly higher. The positive correlation between granularity and performance can be explained by the increased levels of activation sparsity in our model, which requires significantly fewer activated neurons (experts). As expected, the performance decreases for the extreme choice of expert size equal to 1 due to significantly higher routing costs. We include additional results for expert granularity with GELU activation in Appendix H.
6 Related Work
Mixture-of-Experts.
MoE layers were introduced as an efficient way to further increase the capacity of deep neural networks applied for NLP tasks, initially in LSTM models [38], and later in Transformers [23]. Since then, they have also been applied to computer vision [33, 5]. MoE layers have gained significant popularity primarily due to their excellent scaling properties [8, 3]. Nonetheless, training such models is challenging, primarily because gating decisions must be discrete to ensure sparse expert selection. Various methods of training were proposed, some of which include reinforcement learning [1], weighting the expert output by the probability to allow computation of the gradient of the router [38], or using the Sinkhorn algorithm [3]. Some of those approaches also suffer from the possibility of load imbalance and therefore require auxiliary losses or alternative expert selection methods [9, 54]. Interestingly, in many cases fixed routing functions perform similarly to trainable routers [34], which suggests that current solutions are largely suboptimal. MoE models can also be derived from pre-trained dense models by splitting the model weights into experts and independently training the routers for each layer [52, 57], which avoids most of the problems present in end-to-end training.
Activation sparsity in Transformers.
Li et al. [24] show that ReLU-based Transformer models produce sparse activations in their intermediate representations, an effect that is prevalent across architectures, layers, and modalities. They propose a simple rule for keeping only top-$k$ activations in each MLP layer, which results in a model with comparable performance. Similarly, Mirzadeh et al. [27] demonstrate that ReLU activation function in LLMs encourages ensuing activation sparsity that can be leveraged to skip redundant computations. Tuli and Jha [46] take a step further and design a dedicated Transformer architecture accelerator that also exploits activation sparsity, while Liu et al. [25] proposes to predict activation sparsity structure in LLMs and reduce the model latency by skipping redundant computations. Jaszczur et al. [18] demonstrate that it is possible to train Transformer models from scratch with a fixed level of activation sparsity and obtain similar performance. Finally, a related line of works focuses on sparsity in the attention distributions instead of intermediate representations [4]. None of the above-mentioned methods explore induced activation sparsity as a way to increase computational gains, nor do they address variance of the number of sparse activations on a per-token basis.
@@ 7 Conclusion
We introduce Dense to Dynamic-$k$ Mixture-of-Experts (D2DMoE), a novel approach that induces activation sparsity to improve the efficiency of Transformer-based models by converting their layers to Mixture-of-Experts (MoE). We demonstrate the interplay between the activation sparsity of dense models and the efficiency of converted MoEs. Moreover, we introduce regression-based router training and dynamic-$k$ routing, which enable our method to efficiently utilize the induced sparsity. Finally, we show how dense-to-sparse-MoE conversion approaches can be extended to MHA projections and gated MLPs. Our approach is compatible with the existing Transformer architectures and significantly improves upon existing MoE conversion schemes. Our findings contribute to the ongoing efforts to make Transformer models more efficient and accessible for a wider range of applications, especially in resource-constrained environments.
Limitations and Broader Impact
While D2DMoE displays promising results in reducing the computational cost of inference in Transformer models, a few limitations should be acknowledged. Our proposed sparsity enforcement and router training phases require additional training time. This overhead, while small, must be considered when evaluating the benefits of our approach. Moreover, we demonstrate improved performance over existing approaches on common NLP and CV tasks, but the scope of our experiments is restricted due to limited access to computational resources. Further research is needed to explore its applicability to extremely large models.
Our work focuses primarily on fundamental machine learning research and we do not see any specific risks or ethical issues associated with our method. Nevertheless, we recognize the potential for misuse of machine learning technology and advocate for responsible AI practices to mitigate such risks.
@@ Acknowledgments
Filip Szatkowski is supported by National Centre of Science (NCP, Poland) Grant No. 2022/45/B/ST6/02817. Bartosz Wójcik is supported by National Centre of Science (NCP, Poland) Grant No. 2023/49/N/ST6/02513. Simone Scardapane was partly funded by Sapienza grant RG123188B3EF6A80 (CENTS). This paper has been supported by the Horizon Europe Programme (HORIZON-CL4-2022-HUMAN-02) under the project "ELIAS: European Lighthouse of AI for Sustainability", GA no. 101120237. For the purpose of Open Access, the authors have applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, PCSS, CI TASK, WCSS) for providing computer facilities and support within computational grants no. PLG/2023/016393, PLG/2023/016321, and PLG/2024/017385.
Appendix A Difference between router training in D2DMoE and MoEfication
Our router training procedure is similar to the one proposed in MoEfication [52], but the source code of the method provided by the authors [^3] MoEfication source code for router training is publicly available at: https://github.com/thunlp/MoEfication/blob/c50bb850307a36f8a0add6123f56ba309a156d13/moefication/utils.py#L188-L260 contains a different routing objective than the one reported in the paper. While the paper describes their router training objective as a prediction of the sum of ReLU activation values in each expert, the source code uses prediction labels created from the sum of the activations in the intermediate layer divided by the maximum value in the entire batch and minimize the binary cross-entropy loss. Assuming that $a_{k,j}$ is the activation vector in the hidden layer of expert $j$ for sample $k$, the label generation for their router can be expressed as: